A building that is 150 ft tall casts a shadow of 20 feet long. At the same time a tree casts a shadow of 2 ft. How tall is the tree?
Answer by checkley77(12844)

150/20=X/2 CROSS MULTIPLY.
20X=2*150
20X=300
X=300/20
X=15 THE TREE IS 15 FEET TALL.
PROOF:
150/20=15/2
15/2=15/2